SYNCHROTRON BASED HIGH RESOLUTION FAR-IR SPECTROSCOPY OF 1,1-DICHLOROETHYLENE

\underline{REBECCA A. PEEBLES}, LENA F. ELMUTI, and SEAN A. PEEBLES, Department of Chemistry, Eastern Illinois University, 600 Lincoln Ave., Charleston, IL 61920; DANIEL A. OBENCHAIN, Department of Chemistry, Wesleyan University, 52 Lawn Avenue, Middletown, CT 06459-0180.

Six vibrational bands of the ^{35}Cl_{2}C=CH_{2} isotopologue of 1,1-dichloroethylene have been recorded in the 350 – 1150 cm^{−1} range using the 0.00096 cm^{−1} resolution far-infrared beamline of the Canadian Light Source synchrotron facility. Results from the analysis of one a-type (ν_{9} = 796.01904(8) cm^{−1}, CCl asymmetric stretch) and one c-type (ν_{11} = 868.488626(26) cm^{−1}, CH_{2} flap) band will be presented. Over 6000 transitions have now been fitted for these two bands, with ground state rotational and centrifugal distortion constants fixed to values determined by rotational spectroscopy,^{a} while the upper state constants have been varied. Anharmonic frequency calculations at the MP2/6-311++G(2d,2p) level were instrumental in assigning the dense spectra. Assignment of additional bands around 603 cm^{−1} (b-type, CCl symmetric stretch, ν_{4}) and 456 cm^{−1} (c-type, CCl_{2} flap, ν_{12}), as well as attempts at assigning the mixed ^{35}Cl^{37}Cl isotopologue spectra for ν_{9} and ν_{11}, are in progress.

aZ. Kisiel, L. Pszczółkowski, Z. Naturforsch, \underline{50a}, (1995), 347-351.